# Deterministic Dynamic Edge-Colouring

### v1

@@ 6.1 Structural Corollaries
$|N_{j}|\geq\mathopen{}\left(\frac{\ell}{8\Delta_{\max}^{3}}\right)\mathclose{}^{j}$
provided that for every $i\leq j$, we have no augmenting and consistent $i$-step Vizing chain. In particular, since $N_{j}\subset V(G)$ for all $j$, we conclude that
@@
$\mathopen{}\left(\frac{\ell}{8\Delta_{\max}^{3}}\right)\mathclose{}^{j}\leq|V(G)|\leq n$
and consequently we have:
$i\leq\log_{\frac{\ell}{8\Delta_{\max}^{3}}}n\leq\frac{\log n}{\log\frac{\ell}{8\Delta_{\max}^{3}}}\leq\frac{\log n}{\sqrt{\log n}}=\sqrt{\log n}$
In particular, this means that there must exist some augmenting and consistent Vizing chain with at most $\sqrt{\log n}$ steps and length at most $\ell\cdot{}\sqrt{\log n}=80(\Delta_{\max}+1)^{7}\cdot{}2^{\sqrt{\log n}}\sqrt{\log n}$
At the same time, it follows from Corollary 46 and Lemma 17, that at most $2^{\tilde{O}_{\log\varepsilon}(\sqrt{\log n})}$ of the $i$-step consistent and non-overlapping Vizing chains can reach the Vizing chain. The choice of $\ell=80(\Delta_{\max}+1)^{7}\cdot{}2^{\sqrt{\log n}}$ allows us to set $i\leq\sqrt{\log n}$ while ensuring the existence of augmenting, consistent and non-overlapping $i$-step Vizing chains by Corollary 46. This guarantees the Vizing chain has length at most $2^{\tilde{O}_{\log\varepsilon}(\sqrt{\log n})}$.
At the same time, it follows from Corollary 42 and Lemma 17, that at most $2^{\tilde{O}_{\log\varepsilon}(\sqrt{\log n})}$ of non-overlapping $i$-step processes can reach the Vizing chain. This allows us to exhaustively locate all of these chains if necessary. The combination of these properties is what allows us to build a dynamic data structure for implicitly maintaining a shortest augmenting and consistent Vizing chain for any edge that might be inserted into the dynamic graph. In the next section we will describe the data structures needed to do so.
@@ 6.2 Contents of the data structures
  • $e=(x_{2}x_{1})\in Y_{1}$ with $x_{1}\in X_{1}$ and $x_{2}\in X_{2}$ if $x_{2}$ is an endpoint of the bichromatic path represented by $x_{1}$.
  • $e=(xy)\in Y_{2}$ with $x,y\in X_{2}$ if the nodes in $G$ represented by $x,y$ in $\mathcal{B}(G,c)$ are neighbours via a coloured edge in $G$.
  • $e=(x_{1}x_{2})\in Y_{3}$ with $x_{1}\in X_{1}$ and $x_{2}\in X_{2}$ if $x_{2}$ belongs to the vertex set of the bichromatic path represented by $x_{1}$, and, furthermore, the distance in the bichromatic path, represented by $x_{1}$, from $x_{2}$ to an endpoint of the bichromatic path is at most $\ell$
If $c$ and $\ell$ is clear from the context, we will often leave the arguments out and simply write $\mathcal{B}(G)$. Let $\mathcal{P}(G,c,\ell)$ or simply $\mathcal{P}$, if the arguments are clear from the context, be the set containing all bichromatic paths $P$ that are contained in some maximal bichromatic path $P^{\prime}$ of $G$ under $c$ such that $P$ has length at most $\ell$ and such that $P\cap P^{\prime}$ contains an endpoint of $P^{\prime}$. Then there is a canonical map $r:\mathcal{P}\cup V(G)\rightarrow V(\mathcal{B}(G))$ that maps a vertex in $V(G)$ to its representative in $X_{2}$, and similarly a bichromatic path $P$ to the representative in $X_{1}$ of the unique maximal bichromatic path $P^{\prime}$ containing $P$. Below we state some very basic and straight-forward properties of this type of graphs:
Claim 48.
Let $G$, $\Delta\leq\Delta_{\max}$, $c$, and $\ell$ be as in Definition 47. Then for $\mathcal{B}(G)$ the following holds:
@@
  • The maximum in-degree of $\mathcal{B}(G,c)$ is upper bounded by $2(\Delta_{max}+1)^{2}$.
  • The maximum out-degree of $v\in X_{2}$ is upper bounded by $2(\Delta_{max}+1)^{2}$.
Proof.
It is straight-forward to check that for any $u,v\in V(\mathcal{B}(G))$ at most one edge $(uv)$ can be inserted.
The second item follows from the fact that at most $(\Delta_{\max}+1)^{2}$ bichromatic paths can go through any vertex of $G$. More specifically, if $v\in X_{1}$ then at most $2$ in-edges can be incident to $v$, and if $v\in X_{2}$, then at most $(\Delta_{\max}+1)^{2}$ in-edge of $v$ can belong to $Y_{3}$ and at most $(\Delta_{\max}+1)$ in-edges of $v$ can belong to $Y_{2}$. In this case, no in-edges belong to $Y_{1}$.
Similarly, for the third item, we note that for $v\in X_{2}$ at most $(\Delta_{\max}+1)$ out-edges of $v$ belongs to $Y_{2}$ and at most $(\Delta_{\max}+1)^{2}$ out-edges of $v$ belongs to $Y_{1}$, since $v$ can be the endpoint of at most $(\Delta_{\max}+1)^{2}$ maximal bichromatic paths. No out-edges of $v$ belongs to $Y_{3}$. ∎
Finally, we note that any non-overlapping $i$-step process with step-length at most $\ell$ corresponds to a directed path in $\mathcal{B}(G)$ in a canonical way: simply map $w_{1}+y_{1}+P_{1}+\dots+w_{i}+y_{i}+P_{i}$ to the path $(r(w_{1}),r(y_{1}),r(P_{1}),\dots,r(w_{i}),r(y_{i}),r(P_{i}))$.
Again let $G$, $\Delta\leq\Delta_{\max}$, $c$, and $\ell$ be as in Definition 47. We will maintain the data structures containing the following information:
@@
  • For each edge $(x_{2},x_{1})\in Y_{1}$, we maintain for all $i\leq\sqrt{\log n}$ and for all neighbours $w$ of $y_{1}\in r^{-1}(x_{2})$, the following information in $\mathcal{D}_{i}(e,w)$: either all or the
    $\sqrt{\log n}(\sqrt{\log n}-i)\cdot{}3^{\sqrt{\log n}}(\Delta_{\max}+1)^{3\sqrt{\log n}+2}\cdot{}\ell$
    shortest semi-consistent and augmenting $i$-step processes beginning at $w$ with $y_{1}\in r^{-1}(x_{2})$ and $P_{1}\in r^{-1}(x_{1})$.
In the next section, we will describe how to efficiently maintain the above information under colouring and un-colouring of edges. This is sufficient to easily construct a dynamic algorithm, since the above data structures readily contain information about which edges to recolour in order to extend the maintained colouring.
@@ 6.3 Maintaining the data structures
Maintaining the first item:
In order to explicitly maintain $\mathcal{B}(G)$, we will consider the $(\Delta_{\max}+1)^{2}$ graphs induced in $G$ by any pair of distinct colours $\kappa_{1},\kappa_{2}$. Since the colouring $c$ is at all times proper, these graphs consist only of isolated points, paths, and even cycles. In order to dynamically maintain the representation, we maintain each component in a Top-tree [AHdLT05]. In the case of even cycles, we cannot store the component directly in a dynamic tree data structure, but in that case we store the entire component except for one arbitrary edge in a Top-tree. As ancillary information, we store the number of degree $1$ vertices in the component. Note that there can be only 00, $1$, or $2$ vertices of degree $1$. We use this information to explicitly store $\mathcal{B}(G)$: whenever a maximal bichromatic path is changed, we remove the corresponding node and all its related edges. Next we insert representatives of any un-represented maximal bichromatic paths, and add the edges incident to these representatives as specified by Definition 47.
Maintaining the second item:
@@
Next, we will cleanse the marked data structures and recompute their contents from scratch. We will re-compute the marked data structures using the invariant that all data structures of the form $\mathcal{D}_{i-1}$ should be recomputed before the first data structure of the form $\mathcal{D}_{i}$ is recomputed. Performing the re-computation in the case where $i=1$ is straight-forward as there only is one choice for the $1$-step process. To recompute the structures of the form $\mathcal{D}_{i}$ assuming that all structures of the form $\mathcal{D}_{j}$ with $j\leq i-1$ are correct, now amounts to the following: we perform the first step as prescribed in the description of the data structures. Next we use the data-structures of the form $\mathcal{D}_{i-1}$ along the bichromatic path to find the shortest suggestions of augmenting and consistent $(i-1)$-step processes (here the number of vertices we need to traverse to get to an extension is factored into the suggested length). Then we attempt each extension in increasing size of their lengths. For each attempted extension, we check that it is indeed consistent with our prefix extension by explicitly following the path along and checking the conditions as specified in Definition 15 and the following text. If the extension is not consistent, we don’t include it and attempt the next extension instead. We continue this process until we have enough extensions. It is crucial here that each $i-1$ structure contains enough suggested path that even if the maximum number of them overlap, the data structure can at least suggest the required number of non-overlapping paths. We show that this is indeed the case in the next section.
Performing $\texttt{colour}(e,\kappa)$ is now completely symmetric to the above. Assigning $e$ a colour $\kappa$ also causes some data structures to, potentially, be out-dated. These can now be fixed similarly to before. We provide a more in-depth description below:
Maintaining a bichromatic skeleton:
We will let $F_{\kappa_{1},\kappa_{2}}$ denote the pseudo-forest induced in $G$ by all edges coloured $\kappa_{1}$ or $\kappa_{2}$. For some vertex $v\in G$, we let $PT_{\kappa_{1},\kappa_{2}}(v)$ be the pseudo-tree in $F_{\kappa_{1},\kappa_{2}}$ containing $v$. We store every $PT_{\kappa_{1},\kappa_{2}}$ as a tree $T_{\kappa_{1},\kappa_{2}}$ stored as a Top-tree [AHdLT05] and possibly one extra edge stored in a special slot $S_{\kappa_{1},\kappa_{2}}$. We denote by $T_{\kappa_{1},\kappa_{2}}(v)$ the Top-tree currently containing $v$. Note that we will maintain the invariant that $S_{\kappa_{1},\kappa_{2}}=xy$ only if $x,y$ already belong to the same Top-tree i.e. if $T_{\kappa_{1},\kappa_{2}}(v)=T_{\kappa_{1},\kappa_{2}}(u)$. To perform BCS-colour$(uv,\kappa)$, we do as follows:
  • For all $\kappa^{\prime}\neq\kappa$, we check if $T_{\kappa_{1},\kappa_{2}}(v)=T_{\kappa_{1},\kappa_{2}}(u)$. If this is not the case, then we simply call $\texttt{link}(u,v)$ on the Top-tree data-structures. Otherwise, we store $uv$ in the special slot $S_{\kappa_{1},\kappa_{2}}$.
  • For each Top-tree we store as ancillary information the number of vertices with degree $1$ or 00 contained in each cluster, not counting the boundary vertices. It is straight-forward to maintain this information under join and split.
  • After each modification to the Top-tree, we perform a non-local search for each leaf and check if they are incident to the edge stored in $S_{\kappa_{1},\kappa_{2}}$ (if any such edge). If so, the vertices’ degrees are changed accordingly, and they might no longer be classified as vertices with degree $1$ or 00.
@@
  • To add the incident edges, we access the leaves of the Top-tree to find their neighbours in $G$ as well as the first $\ell$ vertices (from either direction) on the leaf-to-leaf path in the Top-tree. Finally, we add the edges as prescribed by Definition 47.
@@
To perform BCS-uncolour$(uv)$ we only have to modify the first step. The rest of the steps are exactly as before. Instead of calling $\texttt{link}(u,v)$, we call $\texttt{cut}(u,v)$ (or remove $uv$ from $S_{\kappa_{1},\kappa_{2}}$ if it stored there). Finally, if $S_{\kappa_{1},\kappa_{2}}=xy$ is non-empty, we call $\texttt{link}(x,y)$, and set $S_{\kappa_{1},\kappa_{2}}=\emptyset$. After this, we do the same clean-up step as we did in BCS-colour, except that insteado of adding parallel, oppositely directed edges between $r(u)$ and $r(v)$, we remove any such edges.
@@
To perform un-colour$(uv)$ we do as follows:
@@
  • We let $R$ be the set containing representatives of the $1$-neighbourhood of $u$ and $v$ in $G$, all representatives of endpoints of new bichromatic paths of $\mathcal{B}(G)$, and all representatives of former endpoints of no longer-existing bichromatic paths.
@@
  • Having updated all data-structures at level $<i$, we update $\mathcal{D}_{i}(e,w)$ as follows: To update $\mathcal{D}_{i}(x_{1}x_{2},w)$ for $i=1$, construct every augmenting $1$-step process starting at $w_{1}=r^{-1}(w)$ with $y_{1}=r^{-1(x_{2})}$ and $P_{1}\in r^{-1}(x_{1})$. If the process is augmenting, we save the vertices in the process. Note that it is crucial that this information can be efficiently computed without access to any other similar data structures.
    For $i>1$, we ensure that all of the data structures $\mathcal{D}_{j}(e,w)$ have been (re)computed for all $j<i$. To update $\mathcal{D}_{i}(x_{2}x_{1},w)$ we let $w_{1}=r^{-1}(w)$ and $y_{1}=r^{-1}(x_{2})$. We pick a valid choice of $P_{1}\in r^{-1}(x_{1})$ For each valid choice of $P_{1}$, there is a unique extension point $x=x(P_{1})$ corresponding to the last vertex of $P_{1}$ visited, if one traverses $P_{1}$ starting at $y_{1}$. Note that $x$ is among the first $\ell$ vertices of the maximal bichromatic path containing $P_{1}$. For each choice of $P_{1}$ (or equivalently $x$), we do as follows:
@@
      Construct a fan consistent with Lemma 13. We let $P_{2}(x)$ resp. $y_{2}(x)$ be the be bichromatic path resp. the endpoint of $P$ consistent with Lemma 13.
@@
      Next we consider $D_{i-1}(r(y_{2}(x))r(P_{2}(x),x)$ and extract the shortest consistent and augmenting $(i-1)$-step process stored here.
@@
      Then we check whether this extension is consistent with $P_{1}$ according to Definition 15. This is done by simply following the $(i-1)$-step process forward and checking the conditions. If not, then we are happy for now. Otherwise, we take the second smallest extension and repeat the process until we either exhaust all choices or we find a process and concatenate it with $r^{-1}(x)+y_{1}+P_{1}$ to form a consistent $i$-step process. We let $L_{x}$ denote the length of the above $i$-step process.
@@
      We repeat the above step for each choice of $x$, and then store the candidate chains in a BBST.
@@
      Then we extract the shortest chain, store its length and the vertices in the chain in the BBST representing $\mathcal{D}_{i}(x_{2}x_{1},w)$.
    6. 6.
      If the shortest chain went through $x$, then we repeat the above process to find the second shortest chain that does not overlap with $r^{-1}(x)+P_{1}(x)$ as before. Then $L(x)$ is defined in terms of this chain and again inserted into the BBST containing all of the shortest candidate chains.
    7. 7.
      This is repeated until $\mathcal{D}_{i}(x_{2}x_{1},w)$ contains the desired number of chains. If we at one point run out of chains to check, then we also terminate the above step and conclude that $\mathcal{D}_{i}(x_{2}x_{1},w)$ is updated.
    8. 8.
@@
In order to perform colour$(uv,\kappa)$, we do almost the same as above. The only change is that instead of un-colouring $uv$ in $G$, we colour it in $G$ with the colour $\kappa$. Then we update $\mathcal{B}(G)$ accordingly, and then proceed exactly as before.

### v2

@@ 6.1.2 Proof of Proposition
$|N_{i}|=|R_{i}|-|\{w\}\cup\mathopen{}\left(\bigcup\limits_{j=1}^{i-1}|R_{j}|\right)\mathclose{}|\cdot{}4(\Delta_{\max}+1)^{5}\geq\tfrac{\ell|\hat{N}_{i-1}|}{(\Delta_{\max}+1)^{3}}-|\{w\}\cup\mathopen{}\left(\bigcup\limits_{j=1}^{i-1}|R_{j}|\right)\mathclose{}|\cdot{}4(\Delta_{\max}+1)^{5}$
valid extension points for Alice. By the induction hypothesis, we can estimate as follows:
@@
$|\{w\}\cup\mathopen{}\left(\bigcup\limits_{j=1}^{i-1}R_{j}\right)\mathclose{}|\cdot{}4(\Delta_{\max}+1)^{5}$ $\leq 4(1+2(1+\sum\limits_{j=1}^{i-1}|\hat{N}_{j}|))(\Delta_{\max}+1)^{5}$
$\leq 12(1+\sum\limits_{j=1}^{i-1}|\hat{N}_{j}|)(\Delta_{\max}+1)^{5}$
and
$\frac{\ell|\hat{N}_{i-1}|}{(\Delta_{\max}+1)^{3}}\geq\frac{\ell}{2(\Delta_{\max}+1)^{3}}\mathopen{}\left(1+\sum\limits_{j=1}^{i-1}|\hat{N}_{j}|\right)\mathclose{}$
Hence, it follows that if $\ell\geq 10^{4}(10\Delta_{\max}+1)^{50\sqrt{\log n}}$, then we have
$|\{w\}\cup\mathopen{}\left(\bigcup\limits_{j=1}^{i-1}R_{j}\right)\mathclose{}|\cdot{}4(\Delta_{\max}+1)^{5}$ $\leq\tfrac{|R_{i}|}{100}$
and that
$|N_{i}|$ $\geq\mathopen{}\left(\tfrac{\ell}{2(\Delta_{\max}+1)^{3}}-12(\Delta_{\max}+1)^{5}\right)\mathclose{}\mathopen{}\left(1+\sum\limits_{j=1}^{i-1}|\hat{N}_{j}|\right)\mathclose{}$
$\geq\tfrac{\ell}{4(\Delta_{\max}+1)^{3}}\mathopen{}\left(1+\sum\limits_{j=1}^{i-1}|\hat{N}_{j}|\right)\mathclose{}$
The lemma now follows by observing that even if Bob bans $\gamma|N_{i}|$ of these points, we still find:
$|\hat{N}_{i}|\geq\tfrac{99}{100}(1-\gamma)|R_{i}|\geq\tfrac{|R_{i}|}{2}$
and
$|\hat{N}_{i}|\geq\tfrac{\ell}{8(\Delta_{\max}+1)^{3}}\mathopen{}\left(1+\sum\limits_{j=1}^{i-1}|\hat{N}_{j}|\right)\mathclose{}$
Now we show the proposition:
Proof 28 (Proof of Proposition 6.14).
Suppose that Alice has not won after round $\sqrt{\log n}$ concludes. Then by Lemma 6.17 and induction, it follows that $\hat{N}_{\sqrt{\log n}+1}$ has size at least $\mathopen{}\left(\tfrac{\ell}{8(\Delta_{\max}+1)^{3}}\right)\mathclose{}^{\sqrt{\log n}+1}>n$, which is a contradiction as $\hat{N}_{\sqrt{\log n}+1}\subset V(G)$ is not a multi-set.
6.1.3 Theorem
In this subsection, we finally deduce Theorem 6.5, which we restate below for the readers convenience:
Theorem 6.18.
Let $w$ be any node, $y$ any neighbour of $w$ and $P$ any maximal bichromatic path beginning at $y$. Then $w$ is $(a,\sqrt{\log n},\ell,\emptyset,\emptyset,P)$-heavy.
Proof 29.
Suppose for contradiction that $w$ is not $(a,\sqrt{\log n},\ell,\emptyset,\emptyset,P)$-heavy. Then by Lemma 6.8, there exists a graded set $D$ with $D_{i}\leq\ell^{i-1}\cdot{}a$ which $(a,\sqrt{\log n},\ell,\emptyset,\emptyset,P)$-suppresses $w$.
We now claim that Bob has a winning strategy even if the conditions of Proposition 6.14 is met. This would yield the desired contradiction. Hereto, we describe the strategy of Bob. Bob will simply play the strategy $B_{i}=D_{i}$. Since the graded set $D$ $(a,\sqrt{\log n},\ell,\emptyset,\emptyset,P)$-suppresses $w$, it follows that there exist no non-empty $(\sqrt{\log n},\ell,\emptyset,\emptyset,P)$-restricted sets containing strings representing semi-consistent and augmenting stepping processes. As such, if we can show that the strategy is valid with respect to $\gamma$, we achieve the desired contradiction. To this end, observe that by Lemma 6.17 as long as Alice has not won yet, we must have $|N_{i}|\geq(\tfrac{\ell}{8(\Delta_{\max}+1)^{3}})^{i}\geq\ell^{i-1}\tfrac{a}{\gamma}$ for all $i\leq\sqrt{\log n}$, and so Bob’s strategy is indeed valid.
The idea behind the dynamic algorithm and the associated data structures presented in the rest of the section, is to exploit the structural results derived in this section. In particular, we will make Corollary 6.13 algorithmic by dynamically and explicitly maintaining sets of the form outlined in the corollary. The idea is that we can use the approach of Lemma 6.12 to splice together relevant parts of sets of lower order along the bichromatic path to form a higher order set. By carefully repairing and maintaining such sets, we can maintain augmenting stepping processes for all starting points. This allows us to efficiently perform colouring updates, as one can build a Vizing fan around an endpoint of an inserted edge, and then use the stored stepping processes to guide the search for an augmenting Vizing chain. Thus the key part of the algorithm becomes the data structure for maintaining these sets. To do so, we heavily rely on the fact that any colouring or un-colouring of an edge cannot affect too many data structures, as not too many points can actually reach such an edge via semi-consistent stepping processes. Therefore, the data structure can afford to explicitly locate and fix any data structures affected by such a change. We expand upon this in the following subsections, but first we begin by outlining the contents of the data structures that we will maintain.
@@ 6.2 Contents of the data structures
  • $(x_{2}x_{1})\in Y_{1}$ with $x_{1}\in X_{1}$ and $x_{2}\in X_{2}$ if $x_{2}$ is an endpoint of the bichromatic path represented by $x_{1}$.
  • $(xy)\in Y_{2}$ with $x,y\in X_{2}$ if the nodes in $G$ represented by $x,y$ in $\mathcal{B}(G,c)$ are neighbours via a coloured edge in $G$.
  • $(x_{1}x_{2})\in Y_{3}$ with $x_{1}\in X_{1}$ and $x_{2}\in X_{2}$ if $x_{2}$ belongs to the vertex set of the bichromatic path represented by $x_{1}$, and, furthermore, the distance in the bichromatic path, represented by $x_{1}$, from $x_{2}$ to an endpoint of the bichromatic path is at most $\ell-1$.
If $c$ and $\ell$ are clear from the context, we will often leave the arguments out and simply write $\mathcal{B}(G)$. Let $\mathcal{P}(G,c,\ell)$ or simply $\mathcal{P}$, if the arguments are clear from the context, be the set containing all bichromatic paths $P$ that are contained in some maximal bichromatic path $P^{\prime}$ of $G$ under $c$ such that $P$ has length at most $\ell$ and such that $P\cap P^{\prime}$ contains an endpoint of $P^{\prime}$. Then there is a canonical map $r:\mathcal{P}\cup V(G)\rightarrow V(\mathcal{B}(G))$ that maps a vertex in $V(G)$ to its representative in $X_{2}$, and similarly a bichromatic path $P$ to the representative in $X_{1}$ of the unique maximal bichromatic path $P^{\prime}$ containing $P$. Below we state some very basic and straight-forward properties of this type of graphs:
Claim 6.20.
Let $G$, $\Delta\leq\Delta_{\max}$, $c$, and $\ell$ be as in Definition 6.19. Then for $\mathcal{B}(G)$ the following holds:
@@
  • The maximum in-degree of $\mathcal{B}(G,c)$ is upper bounded by $2(\Delta_{\max}+1)^{2}$.
  • The maximum out-degree of $v\in X_{2}$ is upper bounded by $2(\Delta_{\max}+1)^{2}$.
Proof 30.
It is straight-forward to check that for any $u,v\in V(\mathcal{B}(G))$ at most one edge $(uv)$ can be inserted, as $G$ is simple.
The second item follows from the fact that at most $(\Delta_{\max}+1)^{2}$ bichromatic paths can go through any vertex of $G$. More specifically, if $v\in X_{1}$ then at most $2$ in-edges can be incident to $v$, and if $v\in X_{2}$, then at most $(\Delta_{\max}+1)^{2}$ in-edges of $v$ can belong to $Y_{3}$ and at most $(\Delta_{\max}+1)$ in-edges of $v$ can belong to $Y_{2}$. In this case, no in-edges belong to $Y_{1}$.
Similarly, for the third item, we note that for $v\in X_{2}$ at most $(\Delta_{\max}+1)$ out-edges of $v$ belongs to $Y_{2}$ and at most $(\Delta_{\max}+1)^{2}$ out-edges of $v$ belongs to $Y_{1}$, since $v$ can be the endpoint of at most $(\Delta_{\max}+1)^{2}$ maximal bichromatic paths. No out-edges of $v$ belongs to $Y_{3}$.
Finally, we note that any consistent $i$-step process with step-length at most $\ell$ corresponds to a directed path in $\mathcal{B}(G)$ in a canonical way: simply map $w_{1}+y_{1}+P_{1}+\dots+w_{i}+y_{i}+P_{i}$ to the path $(r(w_{1}),r(y_{1}),r(P_{1}),\dots,r(w_{i}),r(y_{i}),r(P_{i}))$. Note that we will never revisit a representative of a path, since different steps are not allowed to be part of the same maximal bichromatic path.
Again let $G$, $\Delta\leq\Delta_{\max}$, $c$, and $\ell$ be as in Definition 6.19. We will maintain data structures containing the following information:
@@
  • For each edge $e=(x_{2},x_{1})\in Y_{1}$, we maintain for all $i\leq\sqrt{\log n}$ and for all neighbours $w$ of $y_{1}\in r^{-1}(x_{2})$, the following information in $\mathcal{D}_{i}(e,w)$: either a non-empty set $\mathcal{S}^{i}[e,w]$ certifying that $w$ is $(\tfrac{a}{2},i,P,y_{1})$-spread (with $P$ maximal bichromatic path containing all paths in $r^{-1}(x_{1})$) or an empty set $\mathcal{S}^{i}[e,w]$. Importantly, we require that if $w$ is $(i,a,\emptyset,\emptyset,x_{1},y_{1})$-heavy, then the set $\mathcal{S}^{i}[e,w]$ is non-empty. In particular, it follows from Theorem 6.5 that all sets of the form $\mathcal{S}^{\sqrt{\log n}}[e,w]$ are non-empty.
In the next section, we will describe how to efficiently maintain the above information under colouring and un-colouring of edges. This is sufficient to construct a dynamic algorithm, since the above data structures readily contain information about which edges to recolour in order to extend the maintained colouring.
@@ 6.3 Maintaining the data structures
In this section, we will consider two main operations: $\texttt{colour}(e,\kappa)$ and $\texttt{un-colour}(e)$. Here $\texttt{colour}(e,\kappa)$ will extend the maintained colouring $c$ by giving the edge $e$ the colour $\kappa$. We require that $\kappa$ is free at both endpoints of $e$ for $\texttt{colour}(e,\kappa)$ to be called. Oppositely, the operation $\texttt{un-colour}(e)$ will reduce the colouring $c$ by assigning $e$ a blank colour $\neg$. In addition to these operations, we will also support $\texttt{add}(e)$ which inserts a new edge $e$ initially coloured $\neg$, and $\texttt{remove}(e)$ which removes a blankly coloured edge $e$ from the graph. These last two operations are at this point straight-forward to maintain, as all the work can be pushed to the two main operations. As mentioned earlier, we will again assume that at all times $\Delta(G)\leq\Delta_{\max}$. We first give an informal overview of the data structures, and then we provide formal description afterwards.
Maintaining the first item:
In order to explicitly maintain $\mathcal{B}(G)$, we will consider the $\Delta_{max+1}\choose 2$} graphs induced in $G$ by any pair of distinct colours $\kappa_{1},\kappa_{2}$. Since the colouring $c$ is at all times proper, these graphs consist only of isolated points, paths, and even cycles. In order to dynamically maintain the representation, we maintain each component in a Top-tree [HolmEtAL]. In the case of even cycles, we cannot store the component directly in a dynamic tree data structure, but in that case we store the entire component except for one arbitrary edge in a Top-tree. As ancillary information, we store the number of degree $1$ vertices in the component. Note that there can be only 0, 11, or $2$ vertices of degree $1$. We use this information to explicitly store $\mathcal{B}(G)$: whenever a maximal bichromatic path is changed, we remove the corresponding node and all its related edges. Next we insert representatives of any un-represented maximal bichromatic paths, and add the edges incident to these representatives as specified by Definition 6.19.
Maintaining the second item:
We first consider the operation $\texttt{un-colour}(e)$. Whenever $e$ is un-coloured some of the information stored in the data structures is no longer valid. In particular, for every non-empty maximal bichromatic path $P$ that previously contained $e$, any $i$-step process that used $P$ might no longer be valid. In a similar vein, some extension close to $e$ could stop being consistent. We can mark every data structure that might be affected by this as follows: any vertex that can reach $P$ via a $j$-step process, could contain faulty information in their data structures for $i\geq j$. We will later bound this number using Claim 6.20 as well as the fact that $e$ belongs to at most $(\Delta_{\max}+1)^{2}$ such $P$’s.
@@
Performing $\texttt{colour}(e,\kappa)$ is now completely symmetric to the above. Assigning $e$ a colour $\kappa$ also causes some data structures to, potentially, be out-dated. These can now be fixed similarly to before.
Next we give a more formal description of the data structures:
Maintaining a bichromatic skeleton:
@@
  • For all $\kappa^{\prime}\neq\kappa$, we check if $T_{\kappa,\kappa^{\prime}}(v)=T_{\kappa,\kappa^{\prime}}(u)$. If this is not the case, then we simply call $\texttt{link}(u,v)$ on the Top-tree data structures. Otherwise, we store $uv$ in the special slot $S_{\kappa,\kappa^{\prime}}$.
  • For each Top-tree we store as ancillary information the number of vertices with degree $1$ or 0 contained in each cluster, not counting the boundary vertices. It is straight-forward to maintain this information under join and split.
  • After each modification to the Top-tree, we perform a non-local search for each leaf and check if they are incident to the edge stored in $S_{\kappa,\kappa^{\prime}}$ (if any such edge). If so, the vertices’ degrees are changed accordingly, and they might no longer be classified as vertices with degree $1$ or 0.
@@
  • To add the incident edges, we access the leaves of the Top-tree to find their neighbours in $G$ as well as the first $\ell$ vertices (from either direction) on the leaf-to-leaf path in the Top-tree. Finally, we add the edges as prescribed by Definition 6.19.
@@
To perform BCS-uncolour(uv)(uv) we only have to modify the first and the last step. The rest of the steps are exactly as before. Instead of calling $\texttt{link}(u,v)$, we call $\texttt{cut}(u,v)$ (or remove $uv$ from $S_{\kappa,\kappa^{\prime}}$ if it stored there). Finally, if $S_{\kappa,\kappa^{\prime}}=xy$ is non-empty, we call $\texttt{link}(x,y)$, and set $S_{\kappa,\kappa^{\prime}}=\emptyset$. After this, we do the same clean-up step as we did in BCS-colour, except that instead of adding parallel, oppositely directed edges between $r(u)$ and $r(v)$, we remove any such edges.
@@
To perform un-colour(uv)(uv) we do as follows:
@@
  • We let $R$ be the set containing representatives of the $2$-hop neighbourhood of $u$ and $v$ in $G$, all representatives of endpoints of new bichromatic paths of $\mathcal{B}(G)$, and all representatives of former endpoints of no longer-existing bichromatic paths.
@@
  • Having updated all data structures at level <i<i, we update $\mathcal{D}_{i}(e,w)$ as follows: To update $\mathcal{D}_{i}(x_{1}x_{2},w)$ for $i=1$, check if the $1$-step process starting at $w_{1}=r^{-1}(w)$ with $y_{1}=r^{-1}(x_{2})$ and $P_{1}\in r^{-1}(x_{1})$ is augmenting. If so, store the set $\mathcal{S}[e,w,i]$ containing this augmenting process. Note that it is crucial that this information can be efficiently computed without access to any other similar data structures.
    For $i>1$, we ensure that all of the data structures $\mathcal{D}_{j}(e,w)$ have been (re)computed for all $j<i$. To update $\mathcal{D}_{i}(x_{2}x_{1},w)$ we let $w_{1}=r^{-1}(w)$ and $y_{1}=r^{-1}(x_{2})$. We proceed as described earlier. If $w_{1}+y_{1}+x_{1}$ has length at most $\ell$, we conclude that it is both short and augmenting, and we simply store this process in $\mathcal{S}[e,w,i]$.
    Otherwise, we consider all valid choices of $P_{1}\in r^{-1}(x_{1})$ of length at most $\ell$. For each valid choice of $P_{1}$, there is a unique extension point $x=x(P_{1})$ corresponding to the last vertex of $P_{1}$ visited, if one traverses $P_{1}$ starting at $y_{1}$. Note that $x$ is among the first $\ell$ vertices of the maximal bichromatic path containing $P_{1}$. For each choice of $P_{1}$ (or equivalently $x$), we do as follows:
@@
      Construct a fan consistent with Lemma 3.5. We let $P_{2}(x)$ resp. $y_{2}(x)$ be the be bichromatic path resp. the endpoint of $P$ consistent with Lemma 3.5.
@@
      Next we check if this extension is blocked in the sense of Lemma 6.12. This is done by simply following every (i1)(i-1)-step process in $\tilde{\mathcal{S}}$ stored in $D_{i-1}((r(y_{2}(x)r(P_{2}(x))),x)$ checking the conditions.
@@
      We repeat the above step for each choice of $x$, noting which extensions are blocked.
@@
      We pick $\tfrac{a}{2}$ arbitrary unblocked choices of $x$ with non empty associated sets $\tilde{\mathcal{S}}$. For each choice, we pick one non-blocked stepping process (represented by a string in the set stored for $x$), say $x+y_{2}+P_{2}+\dots+w_{j}+y_{j}+P_{j}$, increase it to $w+y+P_{1}+x+y_{2}+P_{2}+\dots+w_{j}+y_{j}+P_{j}$, and represent it as a string. We collect all such strings together to form $\mathcal{S}[i,e,w]$.
@@
[…]